# QCD corrections of $e^{+}e^{-}\to J/\psi+c+\bar{c}$ using the principle of maximum conformality

### v1

@@ Abstract
In this paper, we compute the total and differential cross sections for $e^{+}e^{-}\to J/\psi+c+\bar{c}$ at the $B$ factories up to next-to-leading order (NLO) corrections within the framework of nonrelativistic QCD factorization theory. We then obtain an improved pQCD prediction for this process by using the Principle of Maximum Conformality (PMC). We show the first time that the PMC can be applied for any pQCD calculable observable at the total or differential levels via a self-consistent way in perturbation theory. By the recursive use of renormalization group equation, the resultant PMC series satisfies all the requirements of renormalization group, which is free of conventional renormalization scheme-and-scale ambiguities at any fixed-order level. We thus obtain a more precise prompt total cross section at the NLO level, e.g. $\sigma|_{\rm prompt}^{\rm PMC}=0.565^{+0.144}_{-0.125}~{}\text{pb}$. Here the uncertainty is the squared average of those from the $\alpha_{s}$ fixed-point uncertainty $\Delta\alpha_{s}(M_{Z})$, the uncertainty of charm quark mass $\Delta m_{c}$, and an estimated contribution of the uncalculated NNLO-terms as predicted by the Padé approximation approach. The differential cross sections $d\sigma/dP_{J/\psi}$, $d\sigma/d|\cos\theta|$, and $d\sigma/dz$ for $e^{+}e^{-}\to J/\psi+c+\bar{c}$ are further examined. Those results show that by further considering the feed-down contributions, the PMC predictions show better agreement with the Belle measurements.
1 Introduction
@@
Heavy quarkonium production serves as an ideal laboratory for investigating the interplay between the perturbative and nonperturbative effects in Quantum Chromodynamics (QCD) theory. Since the discovery of $J/\psi$ in 1974, it has been a focal point of both theoretical and experimental researches. To characterize quarkonium production, various models have been introduced, including the colour-evaporation model (CEM) Fritzsch:1977ay ; Halzen:1977rs , the color-singlet model (CSM) Chang:1979nn ; Berger:1980ni ; Matsui:1986dk , the nonrelativistic QCD (NRQCD) effective theory Bodwin:1994jh , and etc.. Among these, the NRQCD effective theory offers a systematic approach for distinguishing between short-distance and long-distance effects in quarkonium production. Notable successes have been achieved for interpreting the experimental data on quarkonium production, particularly in the context of unpolarized cross sections for $J/\psi$ hadroproduction, c.f. the reviews Brambilla:2010cs ; Andronic:2015wma ; Lansberg:2019adr ; Chen:2021tmf and references therein. However, there are still challenges to NRQCD. For example, the hadroproduction cross section of $\eta_{c}$ as observed in the LHCb experiments LHCb:2014oii can be effectively explained by the color-singlet contribution alone, suggesting that the color-octet contribution should be very small Butenschoen:2014dra . This observation appears to conflict with the heavy-quark spin symmetry (HQSS) relation that exists between the long-distance matrix elements (LDMEs) of $\eta_{c}$ and $J/\psi$ Han:2014jya ; Zhang:2014ybe ; Belle:2009bxr ; Li:2014fya . Further exploration of processes involving charmonium is essential to test the NRQCD factorization.
@@
A high luminosity $e^{+}e^{-}$ collider has three general features such as cleanliness, democracy, and holism Murayama:1996ec , which are helpful and have some advantages in performing more precise measurements. The inclusive $J/\psi$ production cross section via the $e^{+}e^{-}$ annihilation has been measured by the BaBar collaboration as $2.54\pm 0.21\pm 0.21\mathrm{~{}pb}$ BaBar:2001lfi ; BaBar:2005nic and by the Belle collaboration as $1.45\pm 0.10\pm 0.13\mathrm{~{}pb}$ Belle:2001lqi ; Belle:2002tfa . Many theoretical investigations have been conducted on this production at leading-order (LO) within NRQCD, yielding results for inclusive $J/\psi$ production within the range of $0.6\sim 1.7\mathrm{~{}pb}$ depending on parameter selections Keung:1980ev ; Driesen:1993us ; Yuan:1996ep ; Cho:1996cg ; Schuler:1998az ; Baek:1998yf ; Liu:2002wq ; Hagiwara:2007bq ; Braaten:1995ez ; Wang:2003fw . A subsequent analysis by the Belle collaboration reports Belle:2002tfa
@@
and suggests $\sigma[e^{+}e^{-}\rightarrow J/\psi+X({\mathrm{non}~{}c\bar{c}})]\sim 0.6\mathrm{~{}pb}$. Reference Ioffe:2003gj proposes that distinct $J/\psi$ production mechanisms can be discerned through measurements of $J/\psi$ polarization. Furthermore, a study on the $J/\psi$ polarization in $B\rightarrow J/\psi+X$ employing the Belle detector is presented in Ref. Ichizawa:2000yh .
Regarding the $c\bar{c}$ component, the experimental data issued by the Belle collaboration Belle:2002tfa , $0.87^{+0.21}_{-0.19}\pm 0.17\mathrm{~{}pb}$, is approximately 5 times greater than the LO NRQCD prediction Liu:2002wq . This substantial deviation has been partially reconciled by considering both the next-to-leading-order (NLO) corrections and the feed-down contributions originating from the higher excited states Zhang:2006ay ; Gong:2009ng . Additionally, it has been highlighted in Ref. Nayak:2007mb that the color transfer in associated heavy-quarkonium production may offer significant contributions to this process. The updated Belle experimental measurement Belle:2009bxr provides
@@
A valid perturbative prediction for any physical observable must be independent of any choices of renormalization scale and renormalization scheme. This is the central property of renormalization group invariance (RGI) Stueckelberg:1953dz ; GellMann:1954fq ; Bogoliubov ; Callan:1970yg ; Symanzik:1970rt ; Peterman:1978tb . There is renormalization scale ambiguity inherent in any initial fixed-order perturbative QCD (pQCD) series, resulting in a substantial theoretical uncertainty in pQCD predictions; e.g. it has been found that the initial NLO pQCD cross-sections for $e^{+}e^{-}\to J/\psi+c+\bar{c}$ at the $B$ factories are highly scale dependent. Therefore, it is crucial to eliminate the renormalization scale ambiguity to achieve an accurate fixed-order prediction with maximum precision while satisfying the principle of RGI. In the literature, the principle of maximum conformality (PMC) Brodsky:2011ta ; Brodsky:2011ig ; Brodsky:2012rj ; Mojaza:2012mf ; Brodsky:2013vpa has been proposed to offer a systematic approach for determining the correct magnitude of the strong coupling $\alpha_{s}$ by using the renormalization group equation (RGE). The PMC provides the underlying principle for the well-known BLM method Brodsky:1982gc and extends its procedures unambiguously to all orders. In the limit of small number of colors $N_{c}\to 0$ Brodsky:1997jk , the PMC reduces to the Gell-Mann Low method Gell-Mann:1954yli for QED. The $\{\beta_{i}\}$-terms of the perturbative series determine its $\alpha_{s}$ running behavior by recursively using the RGE. It should be pointed out that two steps need to be done to achieve a correct PMC prediction. The first step is to get all the $\{\beta_{i}\}$-terms of the process, and the second step is to deal with those $\{\beta_{i}\}$-terms properly. That is, only the RGE-involved $\{\beta_{i}\}$-terms should be used to fix the correct magnitude of $\alpha_{s}$, and the $\{\beta_{i}\}$-terms in connection with the quark-mass anomalous dimension $\gamma_{m}$-functions is used to fix the running quark mass Huang:2022rij ; Ma:2024xeq , and etc.. It has been demonstrated that after applying the PMC, the resulting improved perturbative series becomes scale-invariant that is independent of any initial choice of renormalization scale Wu:2018cmb ; Huang:2021hzr . And due to the elimination of the divergent renormalon terms, the convergence of the resultant perturbative series can be naturally improved. Due to the elimination of all the scheme-dependent non-conformal $\{\beta_{i}\}$-terms, the resulting perturbative QCD series transforms into a scheme-independent conformal series; thus the conventional renormalization scheme ambiguity is simultaneously eliminated Wu:2013ei ; Wu:2014iba ; Wu:2019mky ; DiGiustino:2023jiq . In this paper, by applying the PMC scale-setting approach, we will provide a comprehensive and more accurate analysis of the NLO QCD correction to the production channel $e^{+}e^{-}\to J/\psi+c+\bar{c}$ at the $B$ factories with $e^{+}e^{-}$ collision energy $\sqrt{s}=10.6$ GeV.
@@
This remaining parts of the paper are organized as follows. In Sec.II, we give the LO cross section for the process. The calculation of NLO QCD corrections and the PMC procedures are described in Sec.III. Numerical results and discussions are presented in Sec.IV. Section V is reserved for a summary.
2 LO cross section
@@
$iM_{1}$ $=$ $-\frac{-i}{(p_{1}/2+p_{2})^{2}+i\epsilon}\frac{-i}{(p_{1}+p_{2}+p_{3})^{2}+i\epsilon}\cdot\bar{v}_{m_{e}}(k_{2})(ig\gamma^{\mu})u_{m_{e}}(k_{1})\bar{u}_{m_{c}}(p_{2})(-ig_{s}\gamma^{\nu}T^{a})$ (4)
@@
$iM_{2}$ $=$ $-\frac{-i}{(p_{1}/2+p_{2})^{2}+i\epsilon}\frac{-i}{(p_{1}+p_{2}+p_{3})^{2}+i\epsilon}\cdot\bar{v}_{m_{e}}(k_{2})(ig\gamma^{\mu})u_{m_{e}}(k_{1})\bar{u}_{m_{c}}(p_{2})(-ig_{s}\gamma^{\nu}T^{a})$ (5)
@@
$iM_{3}$ $=$ $-\frac{-i}{(p_{1}/2+p_{3})^{2}+i\epsilon}\frac{-i}{(p_{1}+p_{2}+p_{3})^{2}+i\epsilon}\cdot\bar{v}_{m_{e}}(k_{2})(ig\gamma^{\mu})u_{m_{e}}(k_{1})\bar{u}_{m_{c}}(p_{2})\left(-\frac{2}{3}ig\gamma_{\mu}\right)$ (6)
@@
$iM_{4}$ $=$ $-\frac{-i}{(p_{1}/2+p_{3})^{2}+i\epsilon}\frac{-i}{(p_{1}+p_{2}+p_{3})^{2}+i\epsilon}\cdot\bar{v}_{m_{e}}(k_{2})(ig\gamma^{\mu})u_{m_{e}}(k_{1})\bar{u}_{m_{c}}(p_{2})(-ig_{s}\gamma^{\nu}T^{a})$ (7)
@@ 3.1 Virtual NLO correction
Feynman diagrams with a virtual gluon line connected to the quark pair in a meson exhibit Coulomb singularity, which manifests as power divergences in the IR limit of relative momentum. Such singularity can be addressed through $(c\bar{c})[n]$ wave function renormalization Kramer:1995nb ; Gong:2007db . As a result, no Coulomb divergence appears in the resultant pQCD series. In dimensional regularization, there is a simpler way to extract the NRQCD short-distance coefficients using the method of regions Beneke:1997zp . In this method, one can calculate the hard region contributions directly by expanding the relative momentum of the $(c\bar{c})[n]$-pair before carrying out the loop integration. More explicitly, for the present case, one can directly set the relative momentum $q$ to zero before performing loop integration. The Coulomb divergence vanishes during the calculation with dimensional regularization. Moreover, the IR divergences originating from the virtual correction will be cancelled by the IR divergences stemming from the real correction. The UV divergences need to be removed via the process of renormalization. The renormalization scheme is as follows: the on-shell (OS) scheme is employed for the renormalization of the heavy quark field, the heavy quark mass, and the gluon field; the modified minimal subtraction ($\overline{\rm MS}$) scheme is adopted for the renormalization of the strong coupling constant. Using renormalization schemes, the quantities $\delta Z_{i}\equiv Z_{i}-1$ can be derived, e.g.
@@
$\delta Z^{\rm OS}_{3}$ $=$ $\frac{\alpha_{s}}{4\pi}\left[(\beta^{\prime}_{0}-2C_{A})\left(\frac{1}{\epsilon_{\rm UV}}-\frac{1}{\epsilon_{\rm IR}}\right)-\frac{4}{3}T_{F}\left(\frac{1}{\epsilon_{\rm UV}}-\gamma_{E}+{\rm ln}\frac{4\pi\mu_{R}^{2}}{m_{c}^{2}}\right)\right],$
@@ 3.2 Real NLO correction
The real correction exhibits IR divergences, originating from the phase-space integration in the region where the momentum of the final gluon is close to zero. To address these IR divergences in the real correction, we employ the two-cutoff phase-space slicing method Harris:2001sx . Due to the absence of collinear divergence in the current process, we introduce a single cutoff parameter denoted as $\delta_{s}$. Consequently, the phase space for the real correction is partitioned into two distinct regions: the soft region characterized by $E_{4}\leq\sqrt{s}\delta_{s}/2$ and the hard region where $E_{4}>\sqrt{s}\delta_{s}/2$. Here, the gluon energy $E_{4}$ is defined in the center-of-mass (CM) frame of the initial state particles $e^{+}e^{-}$. More explicitly, the real correction can be divided into two parts
@@
$d\sigma_{\rm Real}^{\rm S}$ $=d\sigma_{\rm LO}\left[\frac{C_{F}\alpha_{s}\Gamma(1+\epsilon)}{\pi}\left(\frac{4\pi\mu_{R}^{2}}{s}\right)^{\epsilon}\right]\left\{\left(\frac{1}{\epsilon}-{\rm ln}\,\delta_{s}^{2}\right)\right.\left(1-\frac{\kappa\,p_{2}\cdot p_{3}}{(\kappa^{2}-1)m_{c}^{2}}{\rm ln}\,\kappa^{2}\right)$ (17)
$+\frac{1}{2\beta_{2}}{\rm ln}\left(\frac{1+\beta_{2}}{1-\beta_{2}}\right)+\frac{1}{2\beta_{3}}{\rm ln}\left(\frac{1+\beta_{3}}{1-\beta_{3}}\right)+\frac{2\,\kappa\,p_{2}\cdot p_{3}}{(\kappa^{2}-1)m_{c}^{2}}\left[\frac{1}{4}{\rm ln}^{2}\left(\frac{1-\beta_{2}}{1+\beta_{2}}\right)\right.$
$+{\rm Li}_{2}\left(1-\frac{\,\kappa\,E_{2}(1+\beta_{2})}{v}\right)+{\rm Li}_{2}\left(1-\frac{\,\kappa\,E_{2}(1-\beta_{2})}{v}\right)-\frac{1}{4}{\rm ln}^{2}\left(\frac{1-\beta_{3}}{1+\beta_{3}}\right)$
$-{\rm Li}_{2}\left(1-\frac{E_{3}(1+\beta_{3})}{v}\right)\left.\left.-{\rm Li}_{2}\left(1-\frac{E_{3}(1-\beta_{3})}{v}\right)\right]\right\},$
@@
In the calculations, we employ the package FeynArts Hahn:2000kx to generate Feynman diagrams and amplitudes, and the package FeynCalc Mertig:1990an ; Shtabovenko:2016sxi to perform color and Dirac traces. The packages \$Apart Feng:2012iq and FIRE Smirnov:2008iw are employed for partial fraction and integration-by-parts (IBP) reduction. Following the IBP reduction, all one-loop integrals are reduced to master integrals, which are numerically evaluated using the package LoopTools Hahn:1998yk . The final phase-space integrations are computed with assistance from the package Vegas Lepage:1977sw and its alteration that can be found in the $B_{c}$ meson generator BCVEGPY Chang:2003cq .
3.3 Renormalization scale setting for the NLO total cross section
Combing all the above mentioned contributions, the total cross section of $e^{+}(k_{1})+e^{-}(k_{2})\to\gamma^{*}\to J/\psi(p_{1})+c(p_{2})+\bar{c}(p_{3})$ up to NLO QCD corrections can be expressed as

### v2

@@ Abstract
In this paper, we compute the total and differential cross sections for $e^{+}e^{-}\to J/\psi+c+\bar{c}$ at the $B$ factories up to next-to-leading order (NLO) corrections within the framework of nonrelativistic QCD factorization theory. We then obtain improved pQCD series of those cross sections by using the Principle of Maximum Conformality (PMC). We show that the PMC can be applied for any pQCD calculable observable at the total and differential levels via a self-consistent way in perturbation theory. We observe that a more precise prompt total cross section at the NLO level can be achieved after applying the PMC, e.g. $\sigma|_{\rm prompt}^{\rm PMC}=0.565^{+0.144}_{-0.125}~{}\text{pb}$. Here the uncertainty is the squared average of those from the $\alpha_{s}$ fixed-point uncertainty $\Delta\alpha_{s}(M_{Z})$, the uncertainty of charm quark mass $\Delta m_{c}$, and an estimated contribution of the uncalculated NNLO-terms as predicted by the Padé approximation approach. The differential cross sections $d\sigma/dP_{J/\psi}$, $d\sigma/d|\cos\theta|$, and $d\sigma/dz$ for $e^{+}e^{-}\to J/\psi+c+\bar{c}$ are further examined. Those results show that by further considering the feed-down contributions, the PMC predictions show better agreement with the Belle measurements.
I Introduction
@@
Heavy quarkonium production serves as an ideal laboratory for investigating the interplay between the perturbative and nonperturbative effects in Quantum Chromodynamics (QCD) theory. Since the discovery of $J/\psi$ in 1974, it has been a focal point of both theoretical and experimental researches. To characterize quarkonium production, various models have been introduced, including the colour-evaporation model (CEM) [1, 2], the color-singlet model (CSM) [3, 4, 5], the nonrelativistic QCD (NRQCD) effective theory [6], and etc.. Among these, the NRQCD effective theory offers a systematic approach for distinguishing between short-distance and long-distance effects in quarkonium production. Notable successes have been achieved for interpreting the experimental data on quarkonium production, particularly in the context of unpolarized cross sections for $J/\psi$ hadroproduction, c.f. the reviews [7, 8, 9, 10] and references therein. However, there are still challenges to NRQCD. For example, the hadroproduction cross section of $\eta_{c}$ as observed in the LHCb experiments [11] can be effectively explained by the color-singlet contribution alone, suggesting that the color-octet contribution should be very small [12]. This observation appears to conflict with the heavy-quark spin symmetry (HQSS) relation that exists between the long-distance matrix elements (LDMEs) of $\eta_{c}$ and $J/\psi$ [13, 14, 15, 16]. Further exploration of processes involving charmonium is essential to test the NRQCD factorization.
@@
A high luminosity $e^{+}e^{-}$ collider has three general features such as cleanliness, democracy, and holism [17], which are helpful and have some advantages in performing more precise measurements. The inclusive $J/\psi$ production cross section via the $e^{+}e^{-}$ annihilation has been measured by the BaBar collaboration as $2.54\pm 0.21\pm 0.21\mathrm{~{}pb}$ [18, 19] and by the Belle collaboration as $1.45\pm 0.10\pm 0.13\mathrm{~{}pb}$ [20, 21]. Many theoretical investigations have been conducted on this production at leading-order (LO) within NRQCD, yielding results for inclusive $J/\psi$ production within the range of $0.6\sim 1.7\mathrm{~{}pb}$ depending on parameter selections [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. A subsequent analysis by the Belle collaboration reports [21]
@@
and suggests $\sigma[e^{+}e^{-}\rightarrow J/\psi+X({\mathrm{non}~{}c\bar{c}})]\sim 0.6\mathrm{~{}pb}$. Reference [32] proposes that distinct $J/\psi$ production mechanisms can be discerned through measurements of $J/\psi$ polarization. Furthermore, a study on the $J/\psi$ polarization in $B\rightarrow J/\psi+X$ employing the Belle detector is presented in Ref. [33].
Regarding the $c\bar{c}$ component, the experimental data issued by the Belle collaboration [21], $0.87^{+0.21}_{-0.19}\pm 0.17\mathrm{~{}pb}$, is approximately 5 times greater than the LO NRQCD prediction [28]. This substantial deviation has been partially reconciled by considering both the next-to-leading-order (NLO) corrections and the feed-down contributions originating from the higher excited states [34, 35]. Additionally, it has been highlighted in Ref. [36] that the color transfer in associated heavy-quarkonium production may offer significant contributions to this process. The updated Belle experimental measurement [15] provides
@@
A physical observable should be independent of the choices of renormalization scale and renormalization scheme. This is the central property of renormalization group invariance (RGI) [37, 38, 39, 40, 41, 42]. However there is renormalization scale ambiguity inherent in any initial fixed-order perturbative QCD (pQCD) series due to the mismatching of $\alpha_{s}$ and the coefficients at each perturbative order, resulting in a substantial theoretical uncertainty in pQCD predictions; e.g. it has been found that the initial NLO pQCD cross-sections for $e^{+}e^{-}\to J/\psi+c+\bar{c}$ at the $B$ factories are highly scale dependent. Therefore, it is crucial to eliminate the renormalization scale ambiguity to achieve a more accurate fixed-order prediction.
Many methods have been suggested to achieve the goal, among them the principle of maximum conformality (PMC) [43, 44, 45, 46, 47] has been suggested and developed since 2011. It suggests a systematic approach to achieve a scheme-and-scale invariant conformal series from the initial pQCD series. The PMC provides the underlying principle for the BLM method [48] and extends its procedures from the NLO level to all orders. In the limit of small number of colors $N_{c}\to 0$ [49], the PMC reduces to the Gell-Mann Low method [50] for QED.
The PMC determines the correct magnitude of $\alpha_{s}$ by recursively using the renormalization group equation (RGE). It should be pointed out that two steps need to be done to achieve a correct PMC prediction. The first step is to get all the $\{\beta_{i}\}$-terms of the process, and the second step is to deal with those $\{\beta_{i}\}$-terms properly. That is, only the RGE-involved $\{\beta_{i}\}$-terms should be used to fix the correct magnitude of $\alpha_{s}$, and other $\{\beta_{i}\}$-terms which are connection to the quark-mass anomalous dimension $\gamma_{m}$-functions should be used to fix the running quark mass [51, 52], and etc.. After applying the PMC, the resulting improved perturbative series becomes scale-invariant that is independent of any initial choice of renormalization scale [53, 54]. Due to the elimination of all the scheme-dependent non-conformal $\{\beta_{i}\}$-terms, the resulting perturbative QCD series transforms into a scheme-independent conformal series [55, 56, 57, 58].
In this paper, by applying the PMC, we will provide a comprehensive and more accurate analysis of the NLO QCD correction to the production channel $e^{+}e^{-}\to J/\psi+c+\bar{c}$ at the $B$ factories with $e^{+}e^{-}$ collision energy $\sqrt{s}=10.6$ GeV.
@@
The remaining parts of the paper are organized as follows. In Sec.II, we give the LO cross section for the process. The calculation of NLO QCD corrections and the PMC procedures are described in Sec.III. Numerical results and discussions are presented in Sec.IV. Section V is reserved for a summary.
II LO cross section
@@
$iM_{1}$ $=$ $-\frac{-i}{(p_{1}/2+p_{2})^{2}+i\epsilon}\frac{-i}{(p_{1}+p_{2}+p_{3})^{2}+i\epsilon}$ (4)
$\cdot\bar{v}_{m_{e}}(k_{2})(ig\gamma^{\mu})u_{m_{e}}(k_{1})\bar{u}_{m_{c}}(p_{2})(-ig_{s}\gamma^{\nu}T^{a})$
@@
$\cdot\left(-\frac{2}{3}ig\gamma_{\mu}\right)v_{m_{c}}(p_{3}),$
$iM_{2}$ $=$ $-\frac{-i}{(p_{1}/2+p_{2})^{2}+i\epsilon}\frac{-i}{(p_{1}+p_{2}+p_{3})^{2}+i\epsilon}$ (5)
$\cdot\bar{v}_{m_{e}}(k_{2})(ig\gamma^{\mu})u_{m_{e}}(k_{1})\bar{u}_{m_{c}}(p_{2})(-ig_{s}\gamma^{\nu}T^{a})$
@@
$\cdot(-ig_{s}\gamma_{\nu}T^{a})v_{m_{c}}(p_{3}),$
$iM_{3}$ $=$ $-\frac{-i}{(p_{1}/2+p_{3})^{2}+i\epsilon}\frac{-i}{(p_{1}+p_{2}+p_{3})^{2}+i\epsilon}$ (6)
$\cdot\bar{v}_{m_{e}}(k_{2})(ig\gamma^{\mu})u_{m_{e}}(k_{1})\bar{u}_{m_{c}}(p_{2})\left(-\frac{2}{3}ig\gamma_{\mu}\right)$
@@
$\cdot(-ig_{s}\gamma_{\nu}T^{a})v_{m_{c}}(p_{3}),$
$iM_{4}$ $=$ $-\frac{-i}{(p_{1}/2+p_{3})^{2}+i\epsilon}\frac{-i}{(p_{1}+p_{2}+p_{3})^{2}+i\epsilon}$ (7)
$\cdot\bar{v}_{m_{e}}(k_{2})(ig\gamma^{\mu})u_{m_{e}}(k_{1})\bar{u}_{m_{c}}(p_{2})(-ig_{s}\gamma^{\nu}T^{a})$
@@ III.1 Virtual NLO correction
Feynman diagrams with a virtual gluon line connected to the quark pair in a meson exhibit Coulomb singularity, which manifests as power divergences in the IR limit of relative momentum. Such singularity can be addressed through $(c\bar{c})[n]$ wave function renormalization [61, 62]. As a result, no Coulomb divergence appears in the resultant pQCD series. In dimensional regularization, there is a simpler way to extract the NRQCD short-distance coefficients using the method of regions [63]. In this method, one can calculate the hard region contributions directly by expanding the relative momentum of the $(c\bar{c})[n]$-pair before carrying out the loop integration. More explicitly, for the present case, one can directly set the relative momentum $q$ to zero before performing loop integration. The Coulomb divergence vanishes during the calculation with dimensional regularization. Moreover, the IR divergences originating from the virtual correction will be cancelled by the IR divergences stemming from the real correction. The UV divergences need to be removed via the process of renormalization. The renormalization scheme is as follows: the on-shell (OS) scheme is employed for the renormalization of the heavy quark field, the heavy quark mass, and the gluon field; the modified minimal subtraction ($\overline{\rm MS}$) scheme is adopted for the renormalization of the strong coupling constant. Using renormalization schemes, the quantities $\delta Z_{i}\equiv Z_{i}-1$ can be derived, e.g.
@@
$\delta Z^{\rm OS}_{3}$ $=$ $\frac{\alpha_{s}}{4\pi}\left[(\beta^{\prime}_{0}-2C_{A})\left(\frac{1}{\epsilon_{\rm UV}}-\frac{1}{\epsilon_{\rm IR}}\right)\right.$
$\left.-\frac{4}{3}T_{F}\left(\frac{1}{\epsilon_{\rm UV}}-\gamma_{E}+{\rm ln}\frac{4\pi\mu_{R}^{2}}{m_{c}^{2}}\right)\right],$
@@ III.2 Real NLO correction
The real correction exhibits IR divergences, originating from the phase-space integration in the region where the momentum of the final gluon is close to zero. To address these IR divergences in the real correction, we employ the two-cutoff phase-space slicing method [64]. Due to the absence of collinear divergence in the current process, we introduce a single cutoff parameter denoted as $\delta_{s}$. Consequently, the phase space for the real correction is partitioned into two distinct regions: the soft region characterized by $E_{4}\leq\sqrt{s}\delta_{s}/2$ and the hard region where $E_{4}>\sqrt{s}\delta_{s}/2$. Here, the gluon energy $E_{4}$ is defined in the center-of-mass (CM) frame of the initial state particles $e^{+}e^{-}$. More explicitly, the real correction can be divided into two parts
@@
$d\sigma_{\rm Real}^{\rm S}$
$=d\sigma_{\rm LO}\left[\frac{C_{F}\alpha_{s}\Gamma(1+\epsilon)}{\pi}\left(\frac{4\pi\mu_{R}^{2}}{s}\right)^{\epsilon}\right]\left\{\left(\frac{1}{\epsilon}-{\rm ln}\,\delta_{s}^{2}\right)\right.$
×$~{}\times\left(1-\frac{\kappa\,p_{2}\cdot p_{3}}{(\kappa^{2}-1)m_{c}^{2}}{\rm ln}\,\kappa^{2}\right)+\frac{1}{2\beta_{2}}{\rm ln}\left(\frac{1+\beta_{2}}{1-\beta_{2}}\right)$
$~{}+\frac{1}{2\beta_{3}}{\rm ln}\left(\frac{1+\beta_{3}}{1-\beta_{3}}\right)+\frac{2\,\kappa\,p_{2}\cdot p_{3}}{(\kappa^{2}-1)m_{c}^{2}}\left[\frac{1}{4}{\rm ln}^{2}\left(\frac{1-\beta_{2}}{1+\beta_{2}}\right)\right.$
$~{}+{\rm Li}_{2}\left(1-\frac{\,\kappa\,E_{2}(1+\beta_{2})}{v}\right)+{\rm Li}_{2}\left(1-\frac{\,\kappa\,E_{2}(1-\beta_{2})}{v}\right)$
$~{}-\frac{1}{4}{\rm ln}^{2}\left(\frac{1-\beta_{3}}{1+\beta_{3}}\right)-{\rm Li}_{2}\left(1-\frac{E_{3}(1+\beta_{3})}{v}\right)$
$~{}\left.\left.-{\rm Li}_{2}\left(1-\frac{E_{3}(1-\beta_{3})}{v}\right)\right]\right\},$ (17)
@@
In doing the calculations, we employ the package FeynArts [65] to generate Feynman diagrams and the related amplitudes, and the package FeynCalc [66, 67] to perform color and Dirac traces. The packages \$Apart [68] and FIRE [69] are employed for partial fraction and integration-by-parts (IBP) reduction. Following the IBP reduction, all one-loop integrals are reduced to master integrals, which are numerically evaluated using the package LoopTools [70]. The final phase-space integrations are computed with assistance from the package Vegas [71] and its alteration that can be found in the $B_{c}$ meson generator BCVEGPY [72].
III.3 The NLO cross sections and the PMC
Combing the above mentioned virtual and real contributions, the total cross section of $e^{+}(k_{1})+e^{-}(k_{2})\to\gamma^{*}\to J/\psi(p_{1})+c(p_{2})+\bar{c}(p_{3})$ up to NLO QCD corrections can be expressed as